OpenStudy (anonymous):

differentiate y=ln2x/sqrtx using quotient rule

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)^'=\frac{gf-'fg'}{g^2}\] with \[f(x)=\ln(2x), f'(x)=\frac{1}{x},g(x)=\sqrt{x}, g'(x)=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

plug and chug with the algebra

OpenStudy (anonymous):

when you are done, you should get \[\frac{2-ln(2x)}{2\sqrt{x^3}}\]